OpenStudy (anonymous):

Please help! Suppose a bank loans $4,500 at an interest rate of 5.5% compounded annually. Find the total amount of interest owed on the loan after 6 years.

OpenStudy (anonymous):

The amount owned on the loan after 6 years is exactly $6204.792645 because A=4500(1.055)^6 A=4500(1.37884281) Remember the formula for compound interest is A=P(1+r/100)^t Where A stands for Amount, P stands for Principal/Loan, R stands for Rate and T stands for time